# Word-specific tonal realizations in Mandarin

### v1

@@ 4 General discussion
Our results might be taken to suggest that word types have their own specific pitch contours stored in the mental lexicon. However, this is not what we are claiming. Within the framework of the discriminative lexicon, pitch contours are not stored. In comprehension, pitch contours contribute to predicting a word’s meaning, and in production, a pitch contour is computed on the fly from the intended meaning. Because the DLM makes no attempt to derive words’ forms from underlying forms, and because no abstract representations mediate comprehension or production (unlike the theories of e.g., Cutler and Clifton Jr., , 1999; Levelt et al., , 1999), the model provides a tool for probing the relation between form and meaning at the level of individual tokens. The facts that for comprehension the mapping from a pitch contour to its context-specific embedding is to a large extent linear, and that for production the reverse mapping is completely linear, indicate that there is considerable isomorphy between the form space of Mandarin word tokens’ pitch contours and the semantic space of Mandarin words’ context-specific meanings. This means that form and meaning mirror each other to a much greater extent than is often assumed, especially in frameworks that take as axiomatic that language has a ‘dual articulation’ (Martinet, , 1965) that allocates form and meaning to two unrelated, orthogonal, components of the grammar.
Since the GPT-2 model used to generate our CEs has been pre-trained on a very large amount of data, the CEs partly capture words’ statistical co-occurence profiles. The isomorphy between pitch contours and CEs therefore suggests that the history of collocational usage affects the realization of tone, and not only the properties of the particular context (such as informativity and the properties of preceding and following words) in which a token occurs. From this perspective, the present results dovetail well with conclusions reached by Seyfarth, (2014) and Sóskuthy and Hay, (2017), who argue that tokens of words that are often highly predictable have shorter spoken word duration even in contexts in which the factors normally driving shortening are absent (see also Tang and Shaw, , 2021, for spoken word duration, maximum pitch, and maximum intensity in Mandarin).
The results that we report in the present study not only fit well with recent novel models for understanding lexical processing, but also have important consequences for teaching Mandarin as a second language. If our results generalize from the rise-fall tone pattern to other tone combinations in Mandarin disyllabic words (and ongoing research suggests that indeed this is the case), then presenting second language learners with the typical tones indicated by Pinyin transcriptions can be highly confusing and counterproductive. For instance, a learner presented with the Pinyin of {CJK*}UTF8bsmi學校, xuéxiào ‘school’, but hearing a fall-rise-fall, and noticing the initial descending pitch, will be totally confused about what she is hearing and what (according to the Pinyin) she should be hearing. In fact, for many tokens, the feedback on the tones from the Pinyin is standing in the way of learning how Mandarin speakers actually realize tones. When error-feedback to the learner is itself error-ridden, not much progress can be expected.
@@
To conclude, we have provided a range of observations that are consistent with the possibility that the details of how tones are realized in Taiwan Mandarin disyllabic words is partially determined by meaning in context. If our interpretation of these observations is on the right track, semantics is an important missing player in current phonetic studies of F0 modulation in Mandarin. We believe our empirical findings are sufficiently strong to open up new lines of research on the realization of pitch in tone languages. We believe that our findings will help improve pedagogical methods for teaching Mandarin to L2 learners. Last but not least, we also hope that our findings will contribute to an improved understanding of why deep learning speech processing systems are so remarkably effective and now constitute the state-of-the-art in natural language processing. Our hypothesis is that these systems can pick up systematicities between form and meaning that are not open to human introspection, but that are visible to GAMs and computational models. Crucially, we hypothesize that these systematicities are not only exploited by computational modeling algorithms, but that they are also essential, albeit subliminally, for optimizing human lexical processing in comprehension and production.
Appendix
A. Autocorrelation in pitch contours
Since F0 is a measurement of the frequency with which the vocal folds vibrate, and given the physical constraints on the muscles that control the vocal folds, it is inevitable that F0 changes only slowly with time, and that hence measurements of F0 at successive points in time are autocorrelated. For statistical evaluation, this has a far-reaching consequence: since the observations of pitch at successive points in time are not independent, unavoidably, the residuals of a standard GAM fitted to pitch contours will be also characterized by autocorrelations, violating the assumption of regression modeling that the errors should be independent and uncorrelated. As a consequence, the p-values associated with predictors in a GAM, and the confidence intervals for these predictors, will be too ‘optimistic’.
By way of example, using the current dataset, we fit a simple model predicting log-transformed F0 with speaker gender and normalized time as predictors. As shown in the left-hand panel of Figure A1, the residuals of this model are characterized by substantial autocorrelation. The auto-correlation function indicates that for this model, at lag 1, the correlation of residuals at adjacent time steps is as high as 0.95, and even at a lag of 10 timesteps, the auto-correlation is still substantial, and well above 0.6.
Figure A1: Autocorrelation function plots for the residuals of a GAM model without (left) and with (right) an AR(1) autoregressive process in the residuals.
The mgcv package offers a way of capturing autocorrelations in the residuals by means of building an autoregressive process into the errors. Accordinng to this AR(1)process, the errors are generated as follows:
$\varepsilon_{t}=\rho\varepsilon_{t-1}+\epsilon_{i},\ \ \epsilon_{i}\sim{\cal N}(0,\sigma).$
The residual at time $t$ is modeled as the sum of a proportion $\rho$ of the residual at time $t-1$ and Gaussian noise. For the present data, inclusion of an AR(1) process with $\rho=0.95$ successfully removes most of the autocorrelation in the errors, as shown in the right-hand panel of Figure A1.
Unfortunately, this solution comes with two problems of its own. First, the autocorrelation at lag 1 varies substantially between tokens. As a consequence, a single value of $\rho$ provides only a rough approximation that is too low for some tokens, and too high for many others. This is illustrated in Figure A2. The horizontal axis displays the distribution of by-token autocorrelation values in the errors, calculated based on the residuals of a GAM model without an AR(1) process in the errors. A large majority of tokens have high autocorrelations (median = 0.79), but there are still quite a number of tokens with much lower autocorrelations. The vertical dimension plots the autocorrelation in the residuals from a model with an AR(1) process in the errors (with $\rho$ f= 0.95). Although overall the autocorrelation in the errors is much lower (median = 0.29), there are still tokens with relatively high autocorrelation, and more tokens with a negative autocorrelation. Inter-token variability remains substantial. This demonstrates that the $\rho$ parameter is set to high for some tokens, but too low for others. In other words, the problem of autocorrelated residuals cannot be properly addressed until by-token specification for the autocorrelation parameter is made possible in GAM.
Figure A2: Scatter plot of by-token autocorrelation in the residuals from a model without AR1 incorporated (horizontal) and one with AR1 (vertical). Distributions of autocorrelation are shown by the boxplots.
Second, inclusion of an AR(1) process in the errors gives rise to residuals that do not approximate a normal distribution. For illustration, the Q-Q plots in the left and middle panels of Figure A3 display the distribution of the residuals of the models without and with an AR(1) process respectively. As can be seen, when an AR(1) process is not incorporated (left panel), residuals are already not normally distributed. They come with thick tails at both ends, indicative of being more similar to a $t$ distribution. However, after applying the AR(1) process (right panel), the residuals exhibit a very spiky distribution, with large numbers of small errors hovering around zero, in combination with thin but very long tails. These tails resist adjustment towards normality using a scaled t-distribution (see van Rij et al., , 2019, for a study where the combination of an AR(1) process and a scaled t-distribution was successful).
Figure A3: Q-Q plots for the residuals of a GAM model without (left) and with (right) the AR(1) process.
The large errors in the long tails of the error distribution can be traced back in part to tokens with discontinuous contours. For example, compare panel (a) of Figure A4 with panels (b)-(d). While the former shows a continuous smooth contour, the remaining three exhibit discontinuities. For panels (b)-(d), the break in the middle is due to the presence of a voiceless consonant in the onset of the second syllable. In the case of the contours in panels (c) and (d), creaky voice towards the end of the falling tone gives rise to an abrupt decrease in F0. In principle, one could start new time series after a clear discontinuity, but this raises new questions. What should counts as a ‘clear’ discontinuity? Does it make sense to start a new time series for timepoints with creaky voice? Are the pitch contours during stop closures planned smoothly but not executed, or are the planned pitch contours truly discontinuous? In the light of these considerations, we decided not to incorporate an AR(1) process in the errors.
Figure A4: Panel (a) shows a complete continuous F0 contour. Panel (b)-(d) are examples of discontinuous contours.
B. Context-sensitive embeddings
Classic models of word embeddings build semantic spaces with count-based techniques. They collect the word occurrence statistics in documents or co-occurrence patterns in context windows (Landauer and Dumais, , 1997; Lund and Burgess, , 1996). As there are usually large numbers of documents and collocating words, the resulting embeddings are large and sparse. Therefore, dimension reduction techniques often reduce the embedding size to approximately 300.
More recent embedding models learn compact and dense representations directly from collocations. For example, Mikolov et al., (2013)’s Word2Vec model trains a single-hidden-layer neural network to predict the word given its surrounding context (continuous bag-of-word) or the surrounding context given the target word (skip-gram). The hidden representations learned by the model are the word embeddings. Different embedding models are proposed and differ in their model architecture or their prediction objectives. For instance, FastText(Bojanowski et al., , 2017) uses sub-word features to construct hidden representations, and can better deal with out-of-vocabulary problems. GLoVe (Pennington et al., , 2014), instead of training with CBOW or skip-gram, directly learns the word embeddings that capture global corpus statistics.
Different from the classic count-based models and the later Word2Vec or GLoVe models, contextualized embeddings (CEs) are the model representations of a deep language model (Bengio et al., , 2000; Melamud et al., , 2016; Raffel et al., , 2020). During the training stage, the language model learns how to encode sentential contexts into CEs, given tasks such as predicting the next token or missing tokens in a sentence. Subsequently, during the inference stage, one can extract the CEs as the representations summarising the word and context in the sentence. Interestingly, although the model is not instructed to learn lexical semantics, nor are they provided with explicit sense inventories, the CEs of the same word type nevertheless reflect their context-specific or sense-level usage information (Peters et al., , 2018; Pavlick, , 2022; Garí Soler and Apidianaki, , 2021).
Two kinds of CE-producing models can be distinguished in terms of whether the model has access to the contexts after the predicting words. In a unidirectional setting, the model is trained with a causal mask and language modeling objective, in which it learns to predict the next word only from the preceding context (e.g. GPT, Radford et al., , 2018). In contrast, a bi-directional model learns through a masked language modeling task, predicting a masked word from its surrounding text (e.g. BERT, Devlin et al., , 2019). Compared to bidirectional models, the embeddings of unidirectional models appear to be more aligned with human cognition in language tasks (Schrimpf et al., , 2021; Goldstein et al., , 2022).
C. LDL and ResLDL
Linear Discriminitive Learning (LDL) is one of the computational engines of the Discriminative Lexicon Model (DLM, Baayen et al., , 2019; Heitmeier et al., 2023c, ). It implements linear mappings between word form and meaning. Mathematically, it is equivalent to multivariate multiple regression in statistics. These linear mappings can be estimated straightforwardly with matrix algebra.
By way of example, let $\bm{C}$ and $\bm{S}$ denote word form and word meaning matrix respectively. The rows of the form matrix $\bm{C}$ contain numeric representations of words forms, and the rows of the semantic matrix $\bm{S}$ contain the corresponding semantic representations. The comprehension mapping $\bm{F}$ can be obtained by solving
$\bm{CF}=\bm{S}.$
In the same vein, the production mapping $\mathbf{G}$ can be obtained by solving
$\bm{SG}=\bm{C}.$
There are different methods of obtaining mappings $\bm{F}$ and $\bm{G}$, which are documented in detail in Heitmeier et al., (2021) and Heitmeier et al., 2023a . We made use of the normal equations of regression, solving the inverse using Choresky decomposition.
The model architecture of ResLDL is shown in Figure A5. The model consists of two routes. The route on the right is a linear mapping that implements LDL. On the left is a nonlinear network. The parallel routes of the linear and nonlinear components are similar to the residual connection model, which was first proposed for computer vision(He et al., , 2016) but is also widely used in later models of natural language processing (Vaswani et al., , 2017). The ResLDL takes the linear transformation matrix estimated by the least squares method as it is. This linear matrix is simply frozen, i.e., the parameters (weights) in the matrix are not updated during training. The tuneable parameters are in the nonlinear route, they are trained to capture the structure in the residuals of the linear model.
The nonlinear route is composed of three feed-forward networks (FFNs), with a sigmoid activation functions in between. The FFN itself is a linear transformation, which contains a set of tunable weights used in matrix multiplication and tunable bias terms, which act like intercepts in linear regression. The nonlinearities are introduced with the sigmoid function which maps input values into the [0, 1] interval. This nonlinear function is crucial because, without it, the successive FFN will collapse into only one linear transformation.
The tuneable parameters of the ResLDL model are initialized randomly, and subsequently optimized during training to minimize the difference between the predictions and the target vectors. At first, the model does a forward pass. For example, in the case of comprehension, the input is a 50-dimensional pitch contour vector, and the model is trying to predict the 768-dimensional CE vector. The model takes a $50\times 768$ LDL comprehension matrix for its linear route and produces a linear prediction. In addition to this linear prediction, the model also takes the pitch contour vector and inputs it into its nonlinear route. This vector is transformed by the first FFN, followed by a sigmoid function, into a latent space, which is set to 200 in this model, resulting in a 200-dimensional latent vector. This latent vector is again transformed by the following FFNs, which ultimately produce a predicted residual vector. The final output vector is the sum of the linear prediction and the predicted residual vector.
The forward pass is followed by a backward pass, using the backpropagation algorithm to update the parameters. First, the output vector is compared to the real vector, and the difference, measured by the mean squared error, is computed. Secondly, the algorithm derives the parameters’ gradients from the computed difference, i.e., the directions with which the parameters should be updated to bring the predictions closer to the real vectors. Finally, the algorithm updates each parameter by subtracting the gradients multiplied by a scalar, the learning rate. The backward pass is complete with the update of all parameters. Training then proceeds with the next forward pass, using the newly updated parameters.
The model has two hyperparameters: the learning rate, and the number of epochs, i.e., the number of times the model parameters are updated on the dataset. If the learning rate is set too low, the optimisation will be too slow, while if it is set too high, we may miss the optimal parameters. Similarly, too few epochs may result in the model being undertrained, and too many iterations may lead to model overfitting. To select optimal hyperparameters, we split the data into three parts: the training set, which contains 3,022 tokens, a validation set (378 tokens), and a test set (378 tokens). The training set is used to learn the model parameters. The validation set is used to determine optimal values for the hyperparameters. Finally, the model performance is evaluated on the test set. In this study, we explored the following set of hyperparameters: two different numbers of epochs (100, 200) and six learning rates ($5\mathrm{e}$-2, $1\mathrm{e}$-2, $5\mathrm{e}$-3, $1\mathrm{e}$-3, $1\mathrm{e}$-4, $1\mathrm{e}$-5). After the hyperparameter selection, the number of epochs was set to 200, and the learning rate was set to $1\mathrm{e}$-5 for the production task and $1\mathrm{e}$-2 for the comprehension task. We used the AdamW optimizer with $\beta_{1}$ set at 0.9 and $\beta_{2}$ set at 0.999, and weight decay set at 0.01. There are 200K tunable parameters in ResLDL. For each model, the training took 20 seconds on a 12th-generation Intel i7 CPU.
Figure A5: Model architecture of residual LDL (ResLDL). The model has two routes transforming the inputs to the predictions (either pitch contours, or contextualized embeddings). In addition to the linear route (right), which uses the standard linear transformation, the model has a second nonlinear route (left) to account for the relations not captured by the linear mapping. The nonlinear route is a three-layer fully connected neural network with sigmoid activation functions. The model adds the output from both routes to make its final predictions. Compared to the standard LDL, the ResLDL adds a nonlinear route to improve the predictions by capturing the regularities left in the residuals of linear projection.

### v2

@@ 4 General discussion
In our DLM comprehension model, the mapping from a pitch contour to its context-specific embedding is to a large extent linear; in the production model, the reverse mapping is completely linear. These facts indicate that there is considerable isomorphy between the form space of Mandarin word tokens’ pitch contours and the semantic space of Mandarin words’ context-specific meanings. In other words, form and meaning mirror each other to a much greater extent than is often assumed, especially in frameworks that take as axiomatic that language has a ‘dual articulation’ (Martinet,, 1965) that allocates form and meaning to two unrelated, orthogonal, components of the grammar. Furthermore, the existence of this isomorphy means that human language users could in principle exploit the associations between form and meaning to optimize comprehension and production.
The question then arises as to whether listeners actually do make use of the distributional-statistical information that is in the speech signal of Mandarin RF words: we think there is a strong possibility that they do so. The pertinent information is present in the speech signal, so speakers are producing tonal realizations that align with word meaning. But this isomorphy cannot be reduced to the consequences of bio-mechanical constraints on the speech production process. Listeners must be learning the distributional statistics of tone and meaning from the speech to which they are exposed. We hasten to note that the learning of the systematicities between pitch and semantics is in all likelihood a completely subliminal process. It is not necessary for learners to be aware of the subtle modulations of pitch contours in relation to equally subtle nuances in meaning. In our conception of the learning process, successful understanding, token by token, will drive low-level learning in the lexical networks, without conscious reflection and effort being required. [^30] For token-by-token incremental lexical learning, see Heitmeier et al., (2023), and for continuous recalibration in vision, see Marsolek, (2008).
@@
To conclude, we have provided a range of observations that are consistent with the possibility that the details of how tones are realized in Taiwan Mandarin disyllabic words is partially determined by meaning in context. If our interpretation of these observations is on the right track, semantics is an important missing player in current phonetic studies of F0 modulation in Mandarin. We believe our empirical findings are sufficiently strong to open up new lines of research on the realization of pitch in tone languages. We also hope that our findings will contribute to an improved understanding of why deep learning speech processing systems are so remarkably effective and now constitute the state-of-the-art in natural language processing. Our hypothesis is that these systems can pick up systematicities between form and meaning that are not open to human introspection, but that are visible to GAMs and computational models. Crucially, we hypothesize that these systematicities are not only exploited by computational modeling algorithms, but are also essential, albeit subliminally, for optimizing human lexical processing in comprehension and production.